2x-3=-11-2x. This is what I have done so far.
2x=2x-3=-11-2x+2x
3/3=13/3

Answer by AnlytcPhil(1806)

2x-3=-11-2x. This is what I have done so far.

2x=2x-3=-11-2x+2x

--------------------------------------------------------


When solving an equation, it is always a mistake to put
two equal signs on the same line. Only put ONE equal 
sign to a line.

Rules:
1. To get rid of a term on one side of an equation, add 
its opposite to BOTH sides of the equation

2.  Get rid of all terms on the LEFT side which DO NOT
contain x, and leave all terms which DO contain x.

3.  Get rid of all terms on the RIGHT side which DO 
contain x, and leave all terms which DO NOT contain x.

4.  When only the unknown letter and its 
numerical coefficient, such as 5x or -3y, are on the 
left side, and the variable does not appear on the
right side, then get rid of the coefficient of x by
dividing both sides by that numerical coefficient.;


-------------------------------------------------------

We are given

          2x - 3 = -11

The 2x contains x and is on the LEFT, so by rule 2,
we leave it there.

The -3 does not contain x and is on the LEFT, so by
rule 2, we get rid of it.  To do this we use rule 1
and add its opposite, which is +3, to both sides. So
we write on the second line:

      2x - 3 + 3 = -11 + 3

Then we simplify. The -3 and the 3 add to 0 so we just
have 2x on the left. We combine the -11 and the 3 on
the right and get -8.  So we write on the third line:  

              2x = -8

Now by rule 4, we must get rid of the 2 coefficient by
dividing both sides by 2, so we write on the fourth
line:

             2x     -8
            覧覧 = 覧覧
              2      2

Now we do some canceling. The 2's cancel on the left.
And the -8 divided by the 2 on the right gives -4

             1       4 
             2x     -8
            覧覧 = 覧覧
              2      2               
              1      1
 
Notice that the negative sign of the -8 did not cancel 
on the right, so it's still there and we get -4 on the 
right, so we write on the last line:

               x = -4

Edwin